For the formation of 3.65g of hcl gas what volume of hydrogen gas and chlorine gas are required at ntp
1
Expert's answer
2017-07-31T11:47:06-0400
M(HCl) = 3.65g
H2 + Cl2 = 2HCl
V(H2) - ?
n(HCl) = n/M = 3.65g/36, 5g/mol = 0,1mol
V(Cl2) - ?
V = n•Vm;
V(H2) = 0,05mol•22,4L/mol = 1,12 L
V(Cl2) = 0,05mol•22,4L/mol = 1,12 L